Direct Materials and Direct Labor Variance Analysis Abbeville Fixture Company manufactures units in a small manufacturing facility. The units are made from brass. Manufacturing has 30 employees. Each employee presently provides 32 hours of labor per week. Information about a production week is as follows: Standard wage per hour $13.8 Standard labor time per unit 15 min. Standard number of lbs. of brass 1.2 lbs. Standard price per lb. of brass $10 Actual price per lb. of brass $10.25 Actual lbs. of brass used during the week 8,158 lbs. Number of units produced during the week 6,600 Actual wage per hour $14.21 Actual hours for the week (30 employees × 32 hours) 960 hrs. Required: a. Determine the standard cost per unit for direct materials and direct labor. Round the cost per unit to two decimal places. Direct materials standard cost per unit $ Direct labor standard cost per unit $ Total standard cost per unit $ b. Determine the direct materials price variance, direct materials quantity variance, and total direct materials cost variance. Round your answers to the nearest whole dollar. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number. Direct Materials Price Variance $ Direct Materials Quantity Variance $ Total Direct Materials Cost Variance $ c. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Round your answers to the nearest whole dollar. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number. Direct Labor Rate Variance $ Direct Labor Time Variance $ Total Direct Labor Cost Variance $
Answer #1

a. Standard cost per unit for direct materials = 1.20 lbs. x $ 10 = $ 12

Standard cost per unit for direct labor = $ 13.80 x 15 / 60 = $ 3.45

b. Direct materials price variance = $ ( 10 - 10.25 ) * 8,158 lbs.= $ 2,039.50 U

Direct materials quantity variance = ( 6,600 x 1.20 - 8,158 ) * $ 10 = $ 2,380 U

Direct materials cost variance = $ 2,039.50 U + $ 2,380 U = $ 4,419.50 U

Direct materials cost variance = 6,600 x $ 12 - 8,158 x $ 10.25 = $ 4,419.50 U

c. Direct labor rate variance = $ ( 13.80 - 14.21) * $ 960 = $ 393.60 U

Direct labor time variance = ( 6,600 x 15 / 60 - 960 ) * $ 13.80 = - $ 9,522 F

Direct labor cost variance = 6,600 * 3.45 - 960 x 14.21 = - $ 9,128.40 F